Application of Job Order Costing

Scanlon Company has a job-order costing system and applies manufacturing overhead cost to products on the basis of machine-hours. The following estimates were used in preparing the predetermined overhead rate for the most recent year:

Machine-hours ...............................

95,000

Manufacturing overhead cost ........

$1,710,000

During the most recent year, a severe recession in the company’s industry caused a buildup of inventory in the company’s warehouses. The company’s cost records revealed the following actual cost and operating data for the year:

Machine-hours .............................................................................

75,000

Manufacturing overhead cost ......................................................

$1,687,500

Amount of applied overhead in inventories at year-end:

Work in process ........................................................................

$337,500

Finished goods ..........................................................................

$253,125

Amount of applied overhead in cost of goods sold ..................

$759,375

Required:

  1. Compute the company's predetermined overhead rate for the year and the amount of underapplied or overapplied overhead for the year.
  1. Determine the difference between net operating income for the year if the underapplied or overapplied overhead is allocated to the appropriate accounts rather than closed directly to Cost of Goods Sold.
Answer #1

Answer a. We will find the company's predetermined overhead rate for the year, by using the following formula:

Where Manufacturing Overhead Cost = $1,710,000 (As given in the question), and

Machine Hours = 95,000 (As given in the question)

Hence Company's predetermined overhead rate for the year = 1710000 / 95000 = $ 18 / Machine Hours

Now, we will find the amount of underapplied or overapplied overhead for the year as mentioned below:

(Note: As the Actual Manufacturing Overhead Cost is higher than the Applied Overhead; the resultant difference is known as Under Applied Overhead)

Answer b. We will allocate the Applied and Under Applied Overhead as mentioned in the following steps:

Step 1: Prepare the following table from the information given in the question:

Step 2: Find the column of Allocation Proportion (%) as mentioned below:

Allocation Proportion = (Applied Overhead in Particular Inventory / Total Applied Overhead) * 100

Hence, the Allocation Proportion for:

Thus, we get the following table:

Step 3: Allocate the Under Applied Overhead in different inventories, in the same proportion as Applied Overheads are allocated.

Thus, we will allocate the Total Under Applied Overhead of $ 337,500.00 (As obtained in the answer a.) in the proportion (ratio) of 25 : 18.75 : 56.25 ; to Work in Progress, Finished Goods and Cost of Goods Sold respectively.

Hence, we will get the following table:

Now, we will determine the difference between net operating income for the year if the underapplied overhead is allocated to the appropriate accounts rather than closed directly to Cost of Goods Sold as mentioned below:

Said Difference Between Net Operating Income = Total Under Applied Overhead - (Under Applied Overhead in Cost of Goods Sold

= $337,500 - $189843.75

= $ 147,656.25
Answer #1

Answer a. We will find the company's predetermined overhead rate for the year, by using the following formula:

Company's predetermined overhead rate for the year = Predetermined Manufacturing Overhead Cost / Predetermined Machine Hours

Where Manufacturing Overhead Cost = $1,710,000 (As given in the question), and

Machine Hours = 95,000 (As given in the question)

Hence Company's predetermined overhead rate for the year = 1710000 / 95000 = $ 18 / Machine Hours

Now, we will find the amount of underapplied or overapplied overhead for the year as mentioned below:

(Note: As the Actual Manufacturing Overhead Cost is higher than the Applied Overhead; the resultant difference is known as Under Applied Overhead)

Answer b. We will allocate the Applied and Under Applied Overhead as mentioned in the following steps:

Step 1: Prepare the following table from the information given in the question:

Step 2: Find the column of Allocation Proportion (%) as mentioned below:

Allocation Proportion = (Applied Overhead in Particular Inventory / Total Applied Overhead) * 100

Hence, the Allocation Proportion for:

Thus, we get the following table:

Step 3: Allocate the Under Applied Overhead in different inventories, in the same proportion as Applied Overheads are allocated.

Thus, we will allocate the Total Under Applied Overhead of $ 337,500.00 (As obtained in the answer a.) in the proportion (ratio) of 25 : 18.75 : 56.25 ; to Work in Progress, Finished Goods and Cost of Goods Sold respectively.

Hence, we will get the following table:

Now, we will determine the difference between net operating income for the year if the underapplied overhead is allocated to the appropriate accounts rather than closed directly to Cost of Goods Sold as mentioned below:

Said Difference Between Net Operating Income = Total Under Applied Overhead - (Under Applied Overhead in Cost of Goods Sold

= $337,500 - $189843.75

= $ 147,656.25